OpenStudy (anonymous):

Find the exact solution of the following quadratic equation by using the Quadratic Formula. -x^2+7x+9=0

OpenStudy (marissalovescats):

Hi! Do you know the quadratic formula? I'd love to help you :)

OpenStudy (anonymous):

Yes ax^2+bx+c

OpenStudy (marissalovescats):

Well that is the form of a quadratic equation. With that you are able to see your a, b and c values But that's not the quadratic formula. This is the formula to solve for your zeros: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

ok so i substitute 2 7 and 9 into the formula

OpenStudy (marissalovescats):

No you use -1 for a 7 for b and 9 for c

OpenStudy (marissalovescats):

So we have \[\frac{ -7 \pm \sqrt{(7^2)-4(-1)(9)} }{ 2(-1) }\]

OpenStudy (marissalovescats):

And you simplify what's under the radical\[\frac{ -7 \pm \sqrt{49+36}}{ -2 }\]

OpenStudy (marissalovescats):

\[\frac{ -7 \pm \sqrt{85} }{ -2 }\] Do you know how to simplify the square root of 85?

OpenStudy (anonymous):

no can you tell me how

OpenStudy (marissalovescats):

Which actually it can't be simplified

OpenStudy (marissalovescats):

So nevermind that, that's our answer

OpenStudy (anonymous):

yea i see my calculator gave me a decimal but thats alot

OpenStudy (marissalovescats):

Yeah well it's not a perfect square and it can't be simplified so our answer is\[\frac{ -7 \pm \sqrt{85} }{ -2 }\]